OpenStudy (anonymous):

Help? find the length and width of a rectangle with area of 50 meters and a maximum perimeter

OpenStudy (anonymous):

Well I know P= 2L + 2W that turns out to: 2(x+y)

OpenStudy (mathmale):

Hola, Vibarguen! Buenos Dias! What's the formula for the AREA of this rectangle?

OpenStudy (mathmale):

Think for a moment. Which of these two expressions do you want to MAXIMIZE? As soon as I'd typed MAXIMIZE, I realized that it's much more common to MINIMIZE the perimeter. Sure that you've copied the problem correctly?

OpenStudy (anonymous):

Morning ;) A=XY yeah mathmale, I know! I've done the minimize before but I don't know how to maximize

OpenStudy (anonymous):

Take a look: A = 64 = xy P = 2(x + y) P = 2x + 128/x dP/dx = 2 - 128/x^2 d^2P/dx^2 = 256/x^3 x^2 = 64 dP/dx = 0 x = 8 Relative min y = 64/x = 8 8 ft x 8 ft

OpenStudy (mathmale):

Are you familiar with the Second Derivative Test?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh and the problem for that was: Find the length and the width of a rectangle that has an area of 64 ft^2 and a minimum perimeter

OpenStudy (mathmale):

You're certainly on the right track, and Iappreciate your typing out your own work. I immediately noticed that you've typed XY=64. Mind checking that against the original problem?

OpenStudy (mathmale):

Oh. A different problem. OK. For all practical purposes, the current problem is very similar to the previous one.

OpenStudy (mathmale):

I notice you found the 2nd derivative and ended up with +256/(x^3). For positive x, that means your second derivative is always positive itself. Positive second derivative tells you what about the critical point? Is that point a max or a min?

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

It tells me it is a min

OpenStudy (mathmale):

Right. My point, exactly. So, if x = Sqrt(50) (in the current problem), your perimeter will be at its minimum (not maximum). So, go back to the formula A=x*y. We know that both x and y must be greater than zero. If you're in doubt, solve A=xy for y and graph the resulting equation. The graph may help you to answer the question posed here.

OpenStudy (anonymous):

so I do the derivative of sqrt 50 ?

OpenStudy (anonymous):

because with this one I am honestly not sure how to do it

OpenStudy (mathmale):

Not quite. Instead, please go back and review what you did in the previous problem. You were right on target there. You found that at x=8, your perimeter will be at its nimimum. My point here is the in the current problem, you are doing exactly the same thing. You have the constraint equation xy=50, or y=50/x. Substitute that into P=2(x+y). Then vary x and determine whether or not the P has a maximum (as opposed to a minimum).

OpenStudy (mathmale):

Hint: P = 2x + 100/x. Try graphing that.

OpenStudy (anonymous):